OpenStudy (anonymous):

find the bound on the magnitude of the error if we approximate sqrt 2 using the taylor approximation of degree three for sqrt 1+x about x=0

OpenStudy (anonymous):

I hate approximations >.<

OpenStudy (anonymous):

Let me see.

OpenStudy (anonymous):

k ty

OpenStudy (anonymous):

i just dont know where to start or what to do

OpenStudy (anonymous):

hey anwara can u plz look at my problem too when u r done here

OpenStudy (anonymous):

I think I am almost there :)

OpenStudy (anonymous):

ty u so much i really appreciate it

OpenStudy (anonymous):

Do you know how to find the Taylor expansion of a function?

OpenStudy (anonymous):

can you please showme

OpenStudy (anonymous):

Hmm, Do you have the formula of finding the Taylor expansion?

OpenStudy (anonymous):

We're just looking for the third degree expansion, that's the first four terms.

OpenStudy (anonymous):

okay my only question is what do i plug into x

OpenStudy (anonymous):

That's good. So, we have the first four terms of the expansion: \[\sqrt{1+x}=1+{x \over 2}-{x^2 \over 8}+{x^3 \over 16}+....\] We are looking for the approximation of sqrt(2), that's sqrt(1+1). So, we will plug x=1, and then we get: \[\sqrt{1+1}=\sqrt{2}=1+{1 \over 2}-{1 \over 8}+{1 \over 16}={23 \over 16}\]

OpenStudy (anonymous):

hey anwara can u plz look at my problem too when u r done here 14 minutes ago

OpenStudy (anonymous):

okay so now what do i do with this value

OpenStudy (anonymous):

Yeah sure rsaad2.

OpenStudy (anonymous):

its on work done

OpenStudy (anonymous):

Now this is the approximated value of sqrt(2), we're looking now for the error in this value.

OpenStudy (anonymous):

Oh I don't really know what the bound on the magnitude of the error exactly is.

OpenStudy (anonymous):

for the taylor series expansion did u use the binomial series to figure it out

OpenStudy (anonymous):

now can u plz look at my problme plz

OpenStudy (anonymous):

Well, I didn't. But, you could, it's actually easier to be found by the binomial series.

OpenStudy (anonymous):

now can u plz look at my problem

OpenStudy (anonymous):

so basically i would plug in sqrt of 1

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

okay so u dont know how do the rest of the problem

OpenStudy (anonymous):

Not really. I don't know what the bound on the magnitude of the error is. If you could provide me with a formula or any information regarding that, I might be able to help.

OpenStudy (anonymous):

if you click on this link you will see it http://math.jasonbhill.com/courses/fall-2010-math-2300-005/lectures/taylor-polynomial-error-bounds

OpenStudy (anonymous):

its under this topic Theorem 10.1 Lagrange Error Bound